# CMT: Mid-Training for Efficient Learning of Consistency, Mean Flow, and Flow Map Models

### v1

@@ 1 Introduction
Figure 1: FID vs. training time for vanilla ECD (Geng et al., 2025b) and CMT (ours) on ImageNet $512{\times}512$. With the proposed mid-training, our CMT w/ ECD (as post-trained flow map) achieves SOTA 2-step FID of 1.84 using only 400 H100 GPU hours (mid- and post-training combined). Under the same budget, vanilla ECD still produces unrecognizable images, and even to reach a reasonable 2-step FID of 3.38 it requires 4643.99 hours. Overall, CMT reduces the total training cost by 91.4% while achieving SOTA.
@@
Theoretically, we show that CMT reduce the gradient discrepancy between the oracle and practical flow map losses, providing a stronger and trajectory-aligned initializer for the flow map post-training. Empirically, we validate our approach on a diverse suite of benchmarks, including pixel space datasets (CIFAR-10, FFHQ $\times64$, AFHQv2 $\times64$, and ImageNet $\times64$) and latent space models for high-resolution synthesis (ImageNet $\times256$ and $\times512$). Initializing flow map models with CMT consistently improves post-training stability, accelerates convergence, and enhances final generation performance. In particular, CMT sets new state-of-the-art (SOTA) 2 step FID scores: 1.97 on CIFAR-10, 1.32 on ImageNet $\times64$, 1.84 on ImageNet $\times512$, 2.34 on AFHQv2 $\times64$, and 2.75 on FFHQ $\times64$. Crucially, these results are achieved with up to a 98% reduction in total training budget and GPU time compared to baselines without our mid-training stage, where training budget is measured by the number of training images processed, equivalently, the number of backpropagated optimization steps. As shown in Figure 1, CMT converge faster, achieving an FID of 1.84 with a 91.4% reduction in training time, compared to the baseline FID of 3.38, on ImageNet $\times512$. On ImageNet $256{\times}256$, CMT attains FID $3.34$ while cutting total training time by $\sim$50% versus MF from scratch (FID $3.43$).
@@
2 Preliminary
2.1 Diffusion Models and Flow Matching.
@@ 2 Preliminary
Relation of the Two Frameworks. The parametrizations are equivalent; the marginal optimal velocity and denoiser satisfy ${\mathbf{v}}({\mathbf{x}}_{t},t)=\Big(\alpha_{t}^{\prime}-\alpha_{t}\tfrac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big){\mathbf{D}}({\mathbf{x}}_{t},t)+\tfrac{\sigma_{t}^{\prime}}{\sigma_{t}}\,{\mathbf{x}}_{t}$, so one can translate between ${\mathbf{v}}_{{\bm{\theta}}}$ and ${\mathbf{D}}_{{\bm{\theta}}}$ given the scheduler. Sampling integrates the PF-ODE (Song et al., 2021), $\frac{\mathrm{d}{\mathbf{x}}_{t}}{\mathrm{d}t}={\mathbf{v}}({\mathbf{x}}_{t},t)$, starting from ${\mathbf{x}}_{T}\sim p_{\mathrm{prior}}$ (Gaussian in both views) down to $t=0$. Either ${\mathbf{v}}_{{\bm{\theta}}}\approx{\mathbf{v}}$ or ${\mathbf{D}}_{{\bm{\theta}}}\approx{\mathbf{D}}$ can be used to realize the drift.
2.2 Few-Step Flow-Map Generative Modeling.
@@
with $d$ a point-wise distance (e.g., squared $\ell_{2}$ or perceptual (Zhang et al., 2018)). At optimum, ${\mathbf{f}}({\mathbf{x}}_{t},t)={\bm{\Psi}}_{t\to 0}({\mathbf{x}}_{t})$ (Proposition F.1). Since ${\bm{\Psi}}_{t\to 0}$ is unavailable, CM uses a stop-gradient surrogate from the previous step,
${\bm{\Psi}}_{t\to 0}({\mathbf{x}}_{t})\;\approx\;{\mathbf{f}}_{{\bm{\theta}}^{-}}\!\left({\mathbf{x}}_{t-\Delta t},\,t-\Delta t\right),\quad\Delta t>0,$
where ${\mathbf{x}}_{t-\Delta t}$ comes from (i) Consistency Distillation (CD): a one-step solver with a pre-trained diffusion teacher, which calls the teacher during training; or (ii) Consistency Training (CT): the analytic estimate ${\mathbf{x}}_{t-\Delta t}=\alpha_{t-\Delta t}{\mathbf{x}}_{0}+\sigma_{t-\Delta t}{\bm{\epsilon}}$ using the same $({\mathbf{x}}_{0},{\bm{\epsilon}})$ as in ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\bm{\epsilon}}$, requiring no teacher calls. Both approaches improve performance by initializing from pre-trained diffusion weights (Lu & Song, 2025; Geng et al., 2025b). In the CT setting, the CM surrogate loss is
@@
$\mathcal{L}_{\mathrm{MF}}({\bm{\theta}}):=\mathbb{E}_{t>s}\mathbb{E}_{{\mathbf{x}}_{t}\sim p_{t}}\Big[w(t)\,\|{\mathbf{h}}_{{\bm{\theta}}}({\mathbf{x}}_{t},t,s)-{\mathbf{h}}_{{\bm{\theta}}^{-}}^{\mathrm{tgt}}({\mathbf{x}}_{t},t,s)\|_{2}^{2}\Big],$
@@
3 Consistency Mid-Training for Efficient and General Flow Map Learning
@@ 3.1 Proposed Pipeline for Flow Map Learning
Stage 1: Pre-Training. Pre-training aims to learn a deterministic ODE sampler that transports samples from $p_{\mathrm{prior}}$ to $p_{\mathrm{data}}$, consistent with the marginals of the forward noising process ${\mathbf{x}}_{t}\sim\mathcal{N}(\alpha_{t}{\mathbf{x}}_{0},\sigma_{t}^{2}{\mathbf{I}})$. A practical choice is an off-the-shelf pre-trained diffusion model with its PF-ODE solver, as many such models are available (Karras et al., 2022; 2024b; Ma et al., 2024; Peebles & Xie, 2023). Alternatively, one may use a lightweight few-step flow map model that supports deterministic sampling (e.g., MF). We refer to these variants collectively as the teacher sampler. Stage 2: Mid-Training (CMT). Efficiently learn a lightweight, trajectory-aligned proxy of the target flow map with minimal computation and stable convergence, without ad-hoc heuristics. CMT’s loss is designed to match the objectives of post-training while using fixed, explicit regression targets supplied by the teacher. Operationally, CMT learns to jump directly between points on the teacher-generated trajectory of a pre-trained model. Because the targets are fixed and high quality, CMT trains stably and yields a trajectory-aligned initializer. Stage 3: Post-Training. Learn the final few-step flow-map model. Compared to random initialization or initialization from pre-trained diffusion models proposed by literature (Geng et al., 2025b; Lu & Song, 2025), the CMT initializer is trajectory-aligned, making post-training more stable, simpler, and faster (as supported by our theoretical analysis in Theorem 5.1 and Appendix F). CMT offers a general recipe for significantly cost-efficient flow map learning.
In what follows, we detail the mid-training stage with CMT, first instantiating it for CM (${\bm{\Psi}}_{t\to 0}$) and then extending it to the general flow map via MF (${\bm{\Psi}}_{t\to s}$).
@@ 3.2 CMT for Learning Consistency Function
We first propose a reinterpretation of $\mathcal{L}_{\mathrm{oracle}\text{-}\mathrm{CM}}$ from a reverse time generative perspective, under which the objective becomes transparent. Every point ${\mathbf{x}}_{t}\sim p_{t}$ along a PF-ODE trajectory is uniquely determined by its terminal state ${\mathbf{x}}_{T}$. Hence, rather than sampling a fresh data point ${\mathbf{x}}_{0}$ and a noise vector $\bm{\epsilon}$} for each $t$, one may sample a single terminal state ${\mathbf{x}}_{T}\sim p_{\mathrm{prior}}$ and trace its entire trajectory backward. Training then reduces to mapping every point on this reverse path to its single consistent origin in the data distribution $p_{\mathrm{data}}$. This yields the following equivalent formulation of the oracle loss; the proof is provided in Section F.1.
@@
If $p_{\mathrm{prior}}$ matches the diffused marginal $p_{T}$ [^1] This holds for sufficiently large $T$ as in EDM, or with appropriate noise schedules as in flow matching., the oracle loss can be expressed as
@@
Building on the reverse-time formulation in Equation 5, we now introduce the training objective of the proposed CMT. We fix a decreasing time discretization $T=t_{M}>t_{M-1}>\cdots>t_{1}>t_{0}=0$. Given a sample ${\mathbf{x}}_{T}\sim p_{\mathrm{prior}}$, we obtain a discrete reference trajectory $\{\hat{{\mathbf{x}}}_{t_{i}}\}_{i=0}^{M}$ by running a numerical ODE solver with the pre–trained diffusion model ${\mathbf{D}}_{\bm{\phi}}$ (in EDM formulation) as the teacher sampler, anchored at $\hat{{\mathbf{x}}}_{t_{M}}={\mathbf{x}}_{T}$. The goal of CMT is for ${\mathbf{f}}_{\bm{\theta}}$ to match any intermediate state $\hat{{\mathbf{x}}}_{t_{i}}$ back to its clean origin $\hat{{\mathbf{x}}}_{t_{0}}$. Training proceeds by minimizing the following loss:
@@ 3.3 CMT  for Learning General Flow Map
We now focus on the MF parameterization for the general flow map learning. MF aims to learn the average drift, defined as
${\mathbf{h}}({\mathbf{x}}_{t},t,s):=\frac{1}{t-s}\int_{s}^{t}{\mathbf{v}}({\mathbf{x}}_{u},u)\mathop{}\!\mathrm{d}u,$
which aggregates the ODE velocity over the interval $[s,t]$.
We observe that this quantity can also be expressed through the flow map. Let ${\mathbf{x}}_{T}$ denote the initial state at time $T$ on the same PF-ODE trajectory as ${\mathbf{x}}_{t}$. Then
@@ 4.1 Experimental Setups
Teachers and Solvers for CMT’s Mid-Training. Across datasets, CMT employs different teacher–solver pairs based on availability: EDM + DPM-Solver++ (Lu et al., 2022; 2025) on {CIFAR10, AFHQv2, FFHQ}; EDM2 + DPM-Solver++ on {ImageNet $\times64$, $\times512$}; and MF-B/4 on ImageNet $\times256$. DPM-Solver++ uses 16 solver steps and MF-B/4 uses 8 with fixed discretization. For mid-training of EDM/EDM2-related settings, we apply a learned perceptual loss to align CMT’s predictions with the teacher’s high-fidelity outputs, specifically using LPIPS (Zhang et al., 2018) in pixel space and ELatentLPIPS (Kang et al., 2024) in latent space. We use squared $\ell^{2}$ loss for MF. We discuss CMT loss function selection motivation in Appendix E.
@@ 4.2 Mid-Training with CMT Improves Flow Map Post-Training
Overall, CMT is both SOTA and highly data-efficient. Notably, CMT with ECT/ECD as the post-training flow map consistently outperforms vanilla ECT/ECD under equal or lower budgets, underscoring the importance of our mid-training. Additional CIFAR-10 evidence appears in Appendix C.1.
@@
MF-XL/2 (Scratch) 0 0 1520 hours 1520 hours 3.43
@@
Remark: Simplicity and Stability of CMT on CM-family Experiments. In mid-training, CMT learns a flow map proxy via an explicit regression target, avoiding stop-gradients, custom time sampling, and handcrafted weights $w(t)$, leading to stable training. Although reference trajectories require a diffusion ODE solver, few-step ($\sim$16) methods such as DPM-Solver++ suffice, and the multistep scheme reuses past states $\hat{\mathbf{x}}_{t_{k}}$ to build later ones, keeping overhead low. On ImageNet $64\times 64$ and $512\times 512$, this yields clear gains: CMT outperforms sCT/sCD while cutting mid- and post-training data and GPU time by 93%–98% (see Appendices B and D).
@@
ImageNet $\times256$: CMT Enables Flexible Teacher Samplers Beyond Diffusion.
@@
Figure 2: FID vs. training time for vanilla MF and CMT (ours) on ImageNet $256{\times}256$. We perform mid-training starting from a randomly-initialized XL/2 model, where CMT of XL/2 size learns to match the deterministic sampler of a weaker, smaller teacher MF-B/4. The resulting mid-trained weights of CMT-XL/2 are then used to initialize MF-XL/2 post-training. This initialization produces semantically meaningful samples early and drives significantly faster convergence. With CMT’s pipeline, training reaches lower FID in only half the GPU hours compared to MF trained from scratch. MF initialized from SiT also converges faster, but requires more than $1520$ hours of pre-training, which exceeds the cost of training MF itself.
@@
We test whether CMT’s mid-training can use a non-diffusion teacher sampler, even if its quality is low. We compare performance on a larger MF-XL/2 model under three settings: (1) MF-XL/2 (scratch): post-train only (vanilla MF with random initialization); (2) MF-XL/2 (SiT init.): pre-train with SiT (Zhu, 2025) followed by post-training with its weights as initialization; (3) CMT-XL/2: train a small MF-B/4 for quick convergence, use it in mid-training as a teacher sampler to generate ODE trajectories for XL/2 with random initialization, then post-train MF-XL/2 initialized from the mid-trained model (see Section B.3).
@@
Table 4 reports the pre-training, mid-training, and post-training time together with the final FID. In particular, CMT-XL/2 cuts total training time by 50% compared to the other two settings, while achieving even better FID. Even though MF-B/4 is a weak teacher (1/2/8-step FID = 24.47/14.96/13.44), using it in CMT substantially accelerates MF-XL/2 training, which converges faster and achieves better FID than vanilla MF. By contrast, SiT pre-training at XL/2 scale requires very long training and its weights lead to unstable MF post-training (Zhu, 2025), making it impractical as a pre-training or mid-training teacher. We also provide qualitative comparisons in Figure 2. After 20 GPU hours, both MF-XL/2 (scratch) and MF-XL/2 (SiT init.) still produce noise, whereas CMT already generates semantically meaningful images. Eventually, CMT attains superior FID with only half the total GPU hours compared to training MF from scratch.

### v2

@@ 1 Introduction
Figure 1: FID vs. training time for vanilla ECD (Geng et al., 2025b) and CMT (ours) on ImageNet $512{\times}512$. With the proposed mid-training, our CMT w/ ECD (as post-trained flow map) achieves SOTA two step FID of 1.84 using only 400 H100 GPU hours (mid- and post-training combined). Under the same budget, vanilla ECD still produces unrecognizable images, and even to reach a reasonable two step FID of 3.38 it requires 4643.99 hours. Overall, CMT reduces the total training cost of flow map models by 91.4% while achieving SOTA performance.
@@
Theoretically, we show that CMT reduce the gradient discrepancy between the oracle and practical flow map losses, providing a stronger and trajectory-aligned initializer for the flow map post-training. Empirically, we evaluate on pixel-space benchmarks (CIFAR-10, FFHQ $64\times 64$, AFHQv2 $64\times 64$, ImageNet $64\times 64$) and latent-space high-resolution models (ImageNet $256\times 256$$512\times 512$) as well as MSCOCO T2I (Lin et al., 2014). Initializing flow map models with CMT improves post-training stability, speeds up convergence, and boosts final quality. CMT achieves new SOTA two-step FIDs of 1.97 (CIFAR-10), 1.32 (ImageNet $64\times 64$), 1.84 (ImageNet $512\times 512$), 2.34 (AFHQv2 $64\times 64$), and 2.75 (FFHQ $64\times 64$), while reducing training budget (images processed, equivalently backprop steps) and GPU time by up to 98% versus baselines without mid-training. On ImageNet $512\times 512$, CMT reaches FID 1.84 with 91.4% less training than the baseline (FID 3.38; Figure 1); on ImageNet $256\times 256$, it attains FID 3.34 with $\sim$50% less training than MF from scratch (FID 3.43). On MSCOCO T2I, CMT achieves the best FID with $\sim$47% less training.
@@
2 Preliminaries and Related Work
2.1 Diffusion Models and Flow Matching
@@
Their Relationship. The parametrizations are equivalent; the marginal optimal velocity and denoiser satisfy ${\mathbf{v}}({\mathbf{x}}_{t},t)=\Big(\alpha_{t}^{\prime}-\alpha_{t}\tfrac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big){\mathbf{D}}({\mathbf{x}}_{t},t)+\tfrac{\sigma_{t}^{\prime}}{\sigma_{t}}\,{\mathbf{x}}_{t}$, so one can translate between ${\mathbf{v}}_{{\bm{\theta}}}$ and ${\mathbf{D}}_{{\bm{\theta}}}$ given the scheduler. Sampling integrates the PF-ODE, $\frac{\mathrm{d}{\mathbf{x}}_{t}}{\mathrm{d}t}={\mathbf{v}}({\mathbf{x}}_{t},t)$, starting from ${\mathbf{x}}_{T}\sim p_{\mathrm{prior}}$ (Gaussian in both views) down to $t=0$. Either ${\mathbf{v}}_{{\bm{\theta}}}\approx{\mathbf{v}}$ or ${\mathbf{D}}_{{\bm{\theta}}}\approx{\mathbf{D}}$ can be used to realize the drift.
2.2 Few-Step Flow Map Generative Modeling
@@
with $d$ a point-wise distance (e.g., squared $\ell_{2}$ or perceptual (Zhang et al., 2018)). At optimum, ${\mathbf{f}}({\mathbf{x}}_{t},t)={\bm{\Psi}}_{t\to 0}({\mathbf{x}}_{t})$ (Proposition F.1). Since ${\bm{\Psi}}_{t\to 0}$ is unavailable, CM uses a stop-gradient surrogate from the previous step: ${\bm{\Psi}}_{t\to 0}({\mathbf{x}}_{t})\;\approx\;{\mathbf{f}}_{{\bm{\theta}}^{-}}\!\left({\mathbf{x}}_{t-\Delta t},\,t-\Delta t\right)$ with $\Delta t>0$, where ${\mathbf{x}}_{t-\Delta t}$ comes from (i) Consistency Distillation (CD): a one-step solver with a pre-trained diffusion teacher, which calls the teacher during training; or (ii) Consistency Training (CT): the analytic estimate ${\mathbf{x}}_{t-\Delta t}=\alpha_{t-\Delta t}{\mathbf{x}}_{0}+\sigma_{t-\Delta t}{\bm{\epsilon}}$ using the same $({\mathbf{x}}_{0},{\bm{\epsilon}})$ as in ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\bm{\epsilon}}$, requiring no teacher calls. Both approaches improve performance by initializing from pre-trained diffusion weights (Lu & Song, 2025; Geng et al., 2025b). In the CT setting, the CM surrogate loss is
@@
$\mathcal{L}_{\mathrm{MF}}({\bm{\theta}}):=\mathbb{E}_{t>s}\mathbb{E}_{{\mathbf{x}}_{t}\sim p_{t}}\Big[w(t)\,\|{\mathbf{h}}_{{\bm{\theta}}}({\mathbf{x}}_{t},t,s)-{\mathbf{h}}_{{\bm{\theta}}^{-}}^{\mathrm{tgt}}({\mathbf{x}}_{t},t,s)\|_{2}^{2}\Big],$ (5)
@@
2.3 Related Work
Diffusion Models and ODE Samplers. The PF-ODE enables deterministic ODE sampling (Song et al., 2021; Lipman et al., 2023; Liu et al., 2023; Karras et al., 2022). DDIM is a first-order (Euler-type) PF-ODE discretization that reduces steps without retraining (Song et al., 2020). DPM-Solver methods use a log-SNR parameterization with 2nd/3rd-order exponential integrators to solve the PF-ODE in very few NFEs (Lu et al., 2022; 2025), and parallel schemes can further optimize the discretization grid (Sabour et al., 2024; Nguyen et al., 2024).
Few-Step Generative Models. Because stepwise ODE integration is slow, distillation approaches train a student to imitate a multi-step teacher via long jumps (Salimans & Ho, 2022; Luhman & Luhman, 2021; Zheng et al., 2023). CM instead learn a direct flow map via consistency property; subsequent work improves training and extends to continuous time (Song et al., 2023; Song & Dhariwal, 2024; Geng et al., 2025b; Lu & Song, 2025). CTM learn maps between arbitrary times but rely on adversarial losses (Kim et al., 2024; Lai et al., 2023). Later work removes the adversarial component with new parameterizations and losses while improving few-step fidelity (Frans et al., 2025; Boffi et al., 2024; Sabour et al., 2025). MF (Geng et al., 2025a) trains from scratch without adversarial loss, at the cost of expensive Jacobian–vector products. Despite recent advances (Lai et al., 2025), training flow map models remains computationally expensive and often unstable.
3 Consistency Mid-Training for Fast, General Flow Map Learning
@@ 3.1 Proposed Pipeline for Flow Map Learning
Stage 1: Pre-Training. It aims to obtain a deterministic ODE sampler that transports samples from $p_{\mathrm{prior}}$ to $p_{\mathrm{data}}$, consistent with the marginals of the forward noising process $\mathcal{N}(\cdot;\alpha_{t}{\mathbf{x}}_{0},\sigma_{t}^{2}{\mathbf{I}})$. A practical choice is an off-the-shelf pre-trained diffusion model with its PF-ODE solver, as many such models are available (Peebles & Xie, 2023; Karras et al., 2024b; Ma et al., 2024). Alternatively, one may use a lightweight few-step flow map model supporting deterministic sampling (e.g., MF). We call these variants collectively the teacher sampler. Stage 2: Mid-Training (CMT). Efficiently learn a lightweight, trajectory-aligned proxy of the target flow map with minimal computation and stable convergence, without ad-hoc heuristics. CMT’s loss is designed to match the objectives of post-training while using fixed, explicit regression targets supplied by the teacher (see Equation 7 or Equation 8 below). Operationally, CMT learns to jump directly between points on the teacher-generated trajectory of a pre-trained model. Because the targets are fixed and high quality, CMT trains stably and yields a trajectory-aligned initializer. Stage 3: Post-Training. Learn the final few-step flow-map model. Compared to random initialization or initialization from pre-trained diffusion models proposed by literature (Geng et al., 2025b; Lu & Song, 2025), the CMT initializer is trajectory-aligned, making post-training more stable, simpler, and faster (as supported by our theoretical analysis in Theorem 5.1 and Appendix F). CMT offers a general recipe for significantly cost-efficient flow map learning.
In what follows, we detail the mid-training stage with CMT, first instantiating it for CM (${\bm{\Psi}}_{t\to 0}$) and then extending it to the general flow map via MF (${\bm{\Psi}}_{t\to s}$), with pseudocode in Algorithm 1.
@@ 3.2 CMT for Learning Consistency Function
We first propose a reinterpretation of $\mathcal{L}_{\mathrm{oracle}\text{-}\mathrm{CM}}$ from a reverse time generative perspective, under which the objective becomes transparent. Every point ${\mathbf{x}}_{t}\sim p_{t}$ along a PF-ODE trajectory is uniquely determined by its terminal state ${\mathbf{x}}_{T}$. Hence, one may sample a single terminal state ${\mathbf{x}}_{T}\sim p_{\mathrm{prior}}$ and trace its entire trajectory backward. Training then reduces to mapping every point on this reverse path to its single consistent origin in the data distribution $p_{\mathrm{data}}$. This yields the following equivalent formulation of the oracle loss; the proof is provided in Section F.1.
@@
If $p_{\mathrm{prior}}$ matches the diffused marginal $p_{T}$ [^1] This holds for sufficiently large $T$ as in EDM, or with appropriate noise schedules as in flow matching. Empirically, $p_{T}$ (data-dependent) and $p_{\mathrm{prior}}$ (data-free) perform identically, so we adopt $p_{\mathrm{prior}}$ in all experiments., the oracle loss can be expressed as
@@
Building on the reverse-time formulation in Equation 6, we now introduce the training objective of the proposed CMT. On the interval $[0,T]$, we fix a time grid $0=t_{0}<t_{1}<\cdots<t_{M}=T$ with $M$ discretization steps. Given a sample ${\mathbf{x}}_{T}\sim p_{\mathrm{prior}}$, we obtain a discrete reference trajectory $\{\hat{{\mathbf{x}}}_{t_{i}}\}_{i=0}^{M}$ by running a numerical ODE solver with the pre–trained diffusion model ${\mathbf{D}}_{\bm{\phi}}$ (in EDM formulation) as the teacher sampler, anchored at $\hat{{\mathbf{x}}}_{t_{M}}={\mathbf{x}}_{T}$. The goal of CMT is for ${\mathbf{f}}_{\bm{\theta}}$ to match any intermediate state $\hat{{\mathbf{x}}}_{t_{i}}$ back to its clean origin $\hat{{\mathbf{x}}}_{t_{0}}$. Training proceeds by minimizing the following loss:
@@ 3.3 CMT  for Learning General Flow Map
We now focus on the MF parameterization for the general flow map learning. MF aims to learn the average drift, defined as ${\mathbf{h}}({\mathbf{x}}_{t},t,s):=\frac{1}{t-s}\int_{s}^{t}{\mathbf{v}}({\mathbf{x}}_{u},u)\mathop{}\!\mathrm{d}u$, which aggregates the ODE velocity over the interval $[s,t]$. We observe that this quantity can also be expressed through the flow map. Let ${\mathbf{x}}_{T}$ denote the initial state at time $T$ on the same PF-ODE trajectory as ${\mathbf{x}}_{t}$. Then
@@ 4.1 Experimental Setups
Teachers and Solvers for CMT’s Mid-Training. Across datasets, CMT employs different teacher–solver pairs based on availability: EDM w/ DPM-Solver++ (Lu et al., 2022; 2025) on {CIFAR10, AFHQv2, FFHQ}; EDM2 w/ DPM-Solver++ on {ImageNet $\times64$, $\times512$}; and MF-B/4 on ImageNet $\times256$. DPM-Solver++ uses 16 solver steps and MF-B/4 uses 8 with fixed discretization. We adopt these solvers with their default log-SNR sampling schedulers due to their strong theoretical foundations and keep all other settings as simple as possible. For mid-training of EDM/EDM2-related settings, we apply a learned perceptual loss to align CMT’s predictions with the teacher’s high-fidelity outputs, specifically using LPIPS (Zhang et al., 2018) in pixel space and ELatentLPIPS (Kang et al., 2024) in latent space. We use squared $\ell^{2}$ loss for MF. We discuss CMT loss function selection motivation in Appendix D.
@@ 4.2 Mid-Training with CMT Improves Flow Map Post-Training
Overall, CMT is both SOTA and highly data-efficient. Notably, CMT with ECT/ECD as the post-training flow map consistently outperforms vanilla ECT/ECD under equal or lower budgets, underscoring the importance of our mid-training. Additional CIFAR-10 evidence appears in Section 4.3.
@@
MF-XL/2 (Rand. Init.) 0 0 1520 hours 1520 hours 3.43
@@
Remark: Simplicity and Stability of CMT on CM-family Experiments. In mid-training, CMT learns a flow map proxy via an explicit regression target, avoiding stop-gradients, custom time sampling, and handcrafted weights $w(t)$, leading to stable training. Although reference trajectories require a diffusion ODE solver, few-step ($\sim$16) methods such as DPM-Solver++ suffice, and the multistep scheme reuses past states $\hat{\mathbf{x}}_{t_{k}}$ to build later ones, keeping overhead low. On ImageNet $64\times 64$ and $512\times 512$, this yields clear gains: CMT outperforms sCT/sCD while cutting mid- and post-training data and GPU time by 93%–98% (see Appendices B and C).
@@
Figure 2: FID vs. training time for vanilla MF and CMT (ours) on ImageNet $256{\times}256$. We perform mid-training starting from a randomly-initialized XL/2 model, where CMT of XL/2 size learns to match the deterministic sampler of a weaker, smaller teacher MF-B/4. The resulting mid-trained weights of CMT-XL/2 are then used to initialize MF-XL/2 post-training. This initialization produces semantically meaningful samples early and drives significantly faster convergence. With CMT’s pipeline, training reaches lower FID in only half the GPU hours compared to MF trained from scratch. MF initialized from SiT also converges fast, but requires more than $1520$ hours of pre-training, which exceeds the cost of training MF itself.
@@
ImageNet $\times256$: CMT Enables Flexible Teacher Samplers Beyond Diffusion. We test whether CMT’s mid-training can use a non-diffusion teacher sampler, even if its quality is low. We compare performance on a larger MF-XL/2 model under three settings: (1) MF-XL/2 (Rand. Init.): post-train only (vanilla MF with random initialization); (2) MF-XL/2 (SiT Init.): pre-train with SiT (Zhu, 2025) followed by post-training with its weights as initialization; (3) CMT-XL/2: train a small MF-B/4 for quick convergence, use it in mid-training as a teacher sampler to generate ODE trajectories for XL/2 with random initialization, then post-train MF-XL/2 initialized from the mid-trained model (see Section B.4).
@@
Table 4 reports the pre-training, mid-training, and post-training time together with the final FID. In particular, CMT-XL/2 cuts total training time by 50% compared to the other two settings, while achieving even better FID. Even though MF-B/4 is a weak teacher (1 / 2 / 8-step FID = 24.47 / 14.96 / 13.44), using it in CMT substantially accelerates MF-XL/2 training, which converges faster and achieves better FID than vanilla MF. By contrast, SiT pre-training at XL/2 scale requires very long training and its weights lead to unstable MF post-training (Zhu, 2025), making it impractical as a pre-training or mid-training teacher. We also provide qualitative comparisons in Figure 2. After 20 GPU hours, both MF-XL/2 (Rand. Init.) and MF-XL/2 (SiT Init.) still produce noise, whereas CMT already generates semantically meaningful images. Eventually, CMT attains superior FID with only half the total GPU hours compared to training MF from scratch.